6. Which gas effuses faster, and by how much: SF6 or Ar
Answer #1

Rate of effusion is inversely proportional to the square root of it's molar mass. Thus, gas with low molar mass will effuse faster. Since Ar has low molecular weight hence it effuse faster.

Rate of Ar/rate of SF6 = √molar mass of Ar/√SF6 molar mass

R(Ar)/R(SF6) = √(M(SF6))/√(M(Ar)) = √146/√40

R(Ar) = 1.91×R(SF6)

So, Ar effuse faster by 1.91 times

Please give feedback!